OpenStudy (anonymous):

a room has a perimeter 48 ft. the width is 2/3 of length. what is the dimensions of the room?

OpenStudy (anonymous):

width = (2/3)x length = x x times (2/3)x = 48 (5/3)x = 48 5x = 144 x = 28.8 therefore length = 28.8 width = 19.2 double check my math but that is the way u solve it

OpenStudy (anonymous):

ya i did something wrong i multilied wrong but same concept

OpenStudy (anonymous):

It should be 2*x + 2*(2/3)x = 48, because we are looking for the perimeter. What you did was find the area instead.10/3*x = 48 -> x = 14.4.